# Quasi-randomization tests for network interference

### v2

@@ 2.3 Null hypothesis of no spillover and conditional randomization test
(Null hypothesis of no spillover effects)
[…]
We highlight that the treatment $G$ is not restricted in the above hypothesis. We now describe the spillover mechanism as a component in the potential outcome function $Y(Z,G)$. In the following example, $Y(Z,G)$ is linearly separable in two components, one of direct effect and another of spillover effect, as stated below.
We re-state the null hypothesis of no spillover effects, as described in Hypothesis 4.
@@
[…]
This approach of defining non-sharp null of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. We can define an exposure function in the above Example 1:
@@
We obtain the setup described in the literature, given a fixed $G$ using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if existing, will be captured in the potential outcome function. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values (Puelz et al. (2022)).
Procedure 0: Consider $C$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|C)$ is imputable under the null.
@@
    Compute p-value as $P_{Z|C}\left(T(Z|C)>T(Z_{obs}|C)\right)$.
The above-stated procedure holds finite-sample validity. The conditioning mechanism $C$ is drawn from $P(C|Z_{obs})$, which the analyst should determine. One must ensure the imputability of the test statistic $T(Z|C)$ when constructing a mechanism for $C$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
@@
$\quad P(Z,G|(Z,G)\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0}),P(Z^{\prime},G^{\prime}|(Z^{\prime},G^{\prime})\in\textbf{Z}\times\textbf{G}\text{ },\mathcal{H}_{0})>0.$
@@ 3 Quasi-randomization test based on network
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, the underlying network structure does not contribute to the total treatment effect under the null of no spillover effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null. This leads to the imputability of a test statistic of choice by fixing the direct treatment assignment $Z$ and randomizing over the network distribution $G$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
@@
Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{obs})$. $T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4.
@@
Note that under $\mathcal{H}_{0}$,
$Y(Z,G)=Y(Z^{\prime},G^{\prime})\qquad Z=Z^{\prime}.$
Thus,
$T_{c}(G|Z_{obs})$ $=T(Z_{obs},G,Y(Z_{obs},G)),$
$=T(Z_{obs},G,Y(Z_{obs},G^{\prime}))\qquad P(G^{\prime}|Z_{obs})>0\text{ under }\mathcal{H}_{0}.$
Since $Z_{obs}$ is arbitrary, we conclude imputability of the $T_{c}(G|Z_{obs})$ under the null $\mathcal{H}_{0}$. ∎
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G)$. An elaborate consideration of the network distribution will be presented in Section 4.
@@
    Consider the randomization distribution $P(G|Z)$.
@@
The proof is given in Appendix B.1. ∎
Note that conditional validity implies unconditional validity of the procedure, too. A crucial component of the proof is identifying the conditional distribution $P(G|Z)$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|C)$. Here, $C$ is a conditioning mechanism consisting of a subset of chosen units. Thus, different experimental designs lead to different $P(Z|C)$. A common problem faced in the literature is implementing the methods in cluster randomized experiments. In a cluster-randomized experiment, the population is classified into clusters where treatments are assigned at the cluster level. If a cluster is treated, all units in that cluster receive the treatment or vice-versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|C)$ degenerate, given the restricted span of cluster treatment assignments space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|C=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of conditional distribution leads to either degeneracy or enormous loss of power. Our methodology captures the null distribution by randomizing over $G|Z$, where the null distribution admits ample support. We discuss this in Sections 3.1.1 and 3.1.2, where we identify the null randomization distribution under different experimental designs.
3.1 Experimental design
Here, we discuss how Procedure 1 can be implemented in different experimental designs. A change in experimental design corresponds to a change in (direct) treatment assignment mechanism. Thus, correctly identifying the distribution $P(G|Z)$ is important to ensure the validity of Procedure 1. We find $P(G|Z)$ for two experimental designs, completely randomized experiments and cluster randomized experiments.
3.1.1 Completely randomized experiment
Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, who are selected in the treatment group. Let $N_{t}$ be fixed. Thus, the remaining units are in the control group. Let $N_{c}=N-N_{t}$ be the size of the units in the control group.
@@ 3.1.1 Completely randomized experiment
We remark that $G\!\perp\!\!\!\perp Z$ by design. Hence, for completely randomized experiments,
$P(G|Z)=P(G).$
Since $P(G)$ is not identified by design, we will explore random graph models in Section 4.1.
3.1.2 Cluster randomized experiment
Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$. That is,
$C_{i}\cap C_{j}=\emptyset$ $\hskip 14.22636pt\forall i\neq j\in[k],$
$\cup_{i=1}^{k}C_{i}=\mathbb{P}.$
Given such a partition, treatment assignment is done at the cluster level instead of the unit level. Following Ugander et al. (2013), we perform a Bernoulli experiment at the cluster level. Thus, we obtain
$W_{i}\sim Bernoulli(p),$
$Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$
[…]
@@ 3.1.2 Cluster randomized experiment
Here, we can see that identifying null randomization distribution becomes challenging in the case of graph cluster randomization experiments. We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
@@ 4 Permutation test via distributional properties
Identification of null randomization distribution $P(G|Z)$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, identifying null randomization distribution becomes challenging in cluster-randomized experiments. This challenge is overcome in our proposed methodology, which is a conditional permutation test. We discuss this in Section 4.1.1.
@@ 4.1 Random graph models
This section describes various random graph models based on which $P(G)$ can be modeled. The Erdős-Rényi random graph model has been studied widely, owing to exact computations of desirable graph properties. We present the model formally in Appendix C.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph better at modeling such real-world networks, proposed by Newman et al. (2002), in Section 4.1.1. For this article, we restrict the models to be covariate-free.
4.1.1 Random graphs with arbitrary degree distribution
@@
Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G)=f(deg(i):i\in\mathbb{P}|p_{1},p_{2},...,p_{N})$. Here, $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution, that is, $p_{i}$ is the probability that a randomly selected unit has degree $i$ and $f(.)$ represents the probability mass function. We call $G$ a random graph with arbitrary degree distribution.
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distribution. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). To deploy Procedure 1, we must identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization but with some uncertainty (Chatterjee et al. (2011)). However, there might be other relational dependencies characterizing the formation of the network that are not captured by random graphs with arbitrary degree distribution. For example, the phenomena of transitivity are not captured in this model, where two nodes have higher chances of being connected if they have a common connecting node.
4.1.2 Exponential random graph model
@@
Common sufficient statistics used in the literature for the characterization of social networks include the number of edges, the number of triangles, and the number of stars of different sizes, say stars of size two or stars of size three (Robins et al. (2007)). It is important to carefully choose the model as an ill-posed ERGM can lead to near degeneracy (Chatterjee and Diaconis (2013); Schweinberger et al. (2020)). As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges given just a single realization.
Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as degree sequence and develop a method conditioning on the sufficient statistic. We then generalize the method to a set of arbitrary sufficient statistics. We use the symmetric nature of the $P(G)$ via an equivalence relation and build a conditional permutation test, as described below.
4.2 Permutation test based on equiprobability events
4.2.1 Test for a random graph with arbitrary degree distribution
Here, we present our main result, which proposes obtaining a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic. We sample the null distribution over all the graphs with the same degree sequence as the observed graph. We obtain a permutation test using the equiprobability of the event set of all the graphs with the same degree sequence. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
@@ 4.2.1 Test for a random graph with arbitrary degree distribution
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most $(N-1)$. Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the graph’s vertices cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
Figure 1: Non-isomorphic graphs with same degree sequence
@@
Consider a random graph $G\in\{0,1\}^{N}\sim P(G)$ where $P(G)=f(deg(i):i\in V|p_{1},p_{2},...,p_{N})$ where $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution of $G$. Let $\mathcal{G}$ be the support of $P(G)$. We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$.
@@
    Define the equivalence class $[G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 6.
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a completely randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a random graph with arbitrary degree distribution. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
@@
Refer to Appendix B.2. ∎
As a corollary, we can substitute any sub-class of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained. For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph in the second step of Procedure 2. We present results for both these Procedures in Section 5.
We emphasize that generating graphs with a given degree sequence admits a polynomial time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by effectively taking many samples to approximate the null distribution.
4.2.2 Test for exponential random graph model
We generalize the above method to a class of exponential random graph models where the sufficient statistics of the model are permutation invariant. We achieve this by conditioning on the graph isomorphism class and generating the null cases from this equivalence class. We formally present the results below.

### v3

@@ 2.3 Null hypothesis of no spillover and conditional randomization test
(Null hypothesis of no spillover effects) $\forall i\in[N]\quad{Y_{i}}_{G_{X_{i}}}(Z,G)={Y_{i}}_{G^{\prime}_{X_{i}}}(Z^{\prime},G^{\prime})$ $\forall(Z,G),(Z^{\prime},G^{\prime})\in\{0,1\}^{N}\times\{0,1\}^{N\times N}\quad,$ $\text{s.t.}\quad Z_{i}=Z_{i}^{\prime},{G_{X}}_{i}={G^{\prime}_{X}}_{i}.$
The treatment $G$ is unrestricted here, fixing the graph covariates. We claim that the null hypothesis, in the above, encodes no spillover effects (see Remark 1). We now describe the spillover mechanism as a component in the potential outcome function $Y(Z,G)$. In the following example, $Y(Z,G)$ is linearly separable into two components, one of direct effect and another of spillover effect, as stated below.
We restate the null hypothesis of no spillover effects, as described in Hypothesis 4.
@@
$\text{s.t.}\quad Z_{i}=Z_{i}^{\prime}.$
This approach of defining non-sharp nulls of spillovers in the literature is through the use of exposure functions (Aronow and Samii (2017); Basse et al. (2019b); Puelz et al. (2022)). The exposure functions are functions of the treatment assignment vector, $Z$, such that restriction sets can be created on $Z$ by partitioning $Z$ on the range of the exposure function. We can define an exposure function in the above Example 1:
@@
We obtain the setup described in the literature, given a fixed $G$, using the above-described exposure function as a restriction set. The setup assumes a spillover mechanism via an exposure function. Since we consider the graph underlying the population to be random, we define restrictions on the joint treatment assignment in Hypothesis 3. A spillover mechanism, if it exists, will be captured in the potential outcome function. We describe the currently proposed method in the literature of a conditional-randomization test procedure to obtain valid p-values using ‘focal units’ (Puelz et al. (2022)).
Procedure 0: Consider $F$ as a subset of units and treatment assignment $Z$ such that a test statistic $T(Z|F)$ is imputable under the null.
@@
    Compute p-value as $P_{Z|F}\left(T(Z|F)>T(Z_{obs}|F)\right)$.
The above-stated procedure holds finite-sample validity. The conditioning mechanism $F$, termed as focal units, is drawn from $P(F|Z_{obs})$, which the analyst should determine. We consider the random selection method as the conditioning mechanism to obtain the focal units. Procedure 0 tests for deviation from the null hypothesis of no spillover effect on focal units. This ensures the imputability of the test statistic $T(Z|F)$. We require the test statistic to be imputable to obtain its sampling distribution under the null, even though we observe only one treatment assignment vector. Below, we define the imputability of a test statistic for the joint treatment assignment.
@@
[…]
@@ 3 Quasi-randomization test based on network
We present the quasi-randomization test for the sharp null of no spillover effects. Our main idea is to generate the null sampling distribution of a test statistic by randomizing over the network-generating process. Intuitively, the underlying network structure also contains information about the spillover effect. This can also be seen in Hypothesis 4 where the network variable $G$ is unrestricted in the, otherwise, non-sharp null keeping the network covariates fixed. Fixing direct treatment assignment and network covariates leads to the imputability of a test statistic of choice. To strictly generalize the conditional randomization test over focal units (Athey et al. (2018)), one can randomize over the network distribution of $G_{F^{c}}$, that is, over the non-focal units in the population. One feature of this generalization is that it ensures the network characteristics of the focal units remain fixed. One can, then, also randomize over the joint treatment assignment vector $(Z,G)$. This expands the state-of-the-art methods over the bivariate treatment assignment vector. The following method focuses on the conditional treatment assignment distribution of $G|Z$. We formally define the conditional test statistic, $T_{c}$, below and show its imputability under the null Hypothesis 4.
@@
Define $T_{c}(G|Z_{obs}):=T(Z_{obs},G,Y_{{G_{X}}_{obs}}(Z_{obs},G))$. $T_{c}(G|Z_{obs})$ is imputable under the null hypothesis of no spillover effects in Hypothesis 4 and is equal to $T(Z_{obs},G,Y_{obs})$. Here, $Y_{obs}=Y_{{G_{X}}_{obs}}(Z_{obs},G_{obs})$.
@@
The proof is given in Supplementary material B.1
We state our proposed quasi-randomization testing procedure below. The procedure assumes the knowledge of the network-generating process $P(G|G_{X})$. An elaborate consideration of the network distribution will be presented in Section 5.
@@
    Consider the randomization distribution $P(G|Z,G_{X})$.
@@
The proof is given in Supplementary material B.3. ∎
Remark 2.
Note that conditional validity implies unconditional validity of the procedure, too. That is, Procedure 1 holds finite sample validity with respect to the bivariate joint distribution of the treatment assignment vector $(Z,G)$, following the law of iterated expectations.
A crucial component of the proof is identifying the conditional distribution $P(G|Z,G_{X})$ for the validity of Procedure 1. This is contingent upon the experimental design, and different experimental designs will lead to different conditional randomization distributions. We compare this with the state-of-the-art methodology presented in Section 2 where the null randomization distribution comes from $P(Z|F)$. Here, $F$ represents focal units. A common problem faced in the literature is the loss of power when implementing the methods in cluster randomized experiments. In a cluster-randomized experiment, the population is classified into clusters where treatments are assigned at the cluster level. If a cluster is treated, all units in that cluster receive the treatment or vice versa. Fixing the treatment assignment of units in the conditioning mechanism might render the null randomization distribution $P(Z|F)$ degenerate, given the restricted span of cluster treatment assignment space. For example, assume there is a focal unit in every cluster. Then, conditioning on the observed treatment of focal units, we obtain $Z|F=\{Z_{obs}\}$ since all the unit treatment assignments can be deterministically imputed by the treatment assignment of the focal unit in that cluster. This reduction in the support of the conditional distribution leads to either degeneracy or enormous loss of power. To get over this, our methodology captures the null distribution by randomizing over the conditional distribution $G|(Z,G_{X})$, where the null distribution admits ample support. We discuss this in Sections 4.1 and 4.2, where we identify the null randomization distribution under different experimental designs.
4 Experimental design
Here, we discuss how Procedure 1 can be implemented in different experimental designs. A change in experimental design corresponds to a change in (direct) treatment assignment mechanism. Thus, correctly identifying the distribution $P(G|Z)$ (and $P(G|Z,G_{X})$) is important to ensure the validity of Procedure 1. We find $P(G|Z)$ for two experimental designs, completely randomized experiments and cluster randomized experiments.
4.1 Completely randomized experiment
Consider a random sample of units of size $N_{t}$ from $\mathbb{P}$, which are selected in the treatment group. Let $N_{t}$ be fixed. Thus, the remaining units are in the control group. Let $N_{c}=N-N_{t}$ be the size of the units in the control group.
@@ 4.1 Completely randomized experiment
We remark that $G\!\perp\!\!\!\perp Z$ by design. Hence, for completely randomized experiments, $P(G|Z)=P(G).$ Since $P(G)$ is not identified by design, we will explore random graph models in Section 5.1.
4.2 Cluster randomized experiment
Consider the following partition of the population $\mathbb{P}$ into clusters $\{C_{1},C_{2},....,C_{k}\}$. That is, $C_{i}\cap C_{j}=\emptyset\forall i\neq j\in[k],\ \ \cup_{i=1}^{k}C_{i}=\mathbb{P}.$ Given such a partition, treatment assignment is done at the cluster level instead of the unit level. Following Ugander et al. (2013), we perform a Bernoulli experiment at the cluster level. Thus, we obtain $W_{i}\sim Bernoulli(p),\ \ Z_{j}=\sum_{i=1}^{k}W_{i}\cdot\mathcal{I}(j\in C_{i}).$ Here, $p$ is the probability that a cluster receives treatment, and the experiment is performed on a given cluster partition of the population. The cluster formation is part of the design of the experiment and is fixed. The experimenter can form these clusters based on the structure of the underlying network. Benchmark graph clustering algorithms can be used to obtain the cluster partition, which depends on graph characteristics such as modularity, cuts, and other related quantities (Newman (2006)). We use $\epsilon$-net clustering to obtain the cluster partition of the network, previously considered in Eckles et al. (2017) and Ugander et al. (2013). We describe the $\epsilon$-net clustering in detail in Supplementary material A. While we present a deterministic version of the clustering procedure, it should be noted that any other algorithm, such as a different selection rule and the resultant $\epsilon$-net cluster output, corresponds to a change in experimental design. Unlike completely randomized experiments, in cluster randomized experiments, $G\not\!\perp\!\!\!\perp Z$. So,
@@ 4.2 Cluster randomized experiment
Here, we can see that identifying a null randomization distribution becomes challenging in the case of graph cluster randomization experiments. We propose a new approach that considers sufficient statistics of the random graph model to build a permutation test.
@@ 5 Permutation test via distributional properties
Identification of null randomization distribution $P(G|Z,G_{X})$ involves knowledge of the distribution of the network-generating process, which is not identified by the design of the experiment. To this end, we explore random graph models to model $P(G)$. As discussed in the previous section, identifying the null randomization distribution becomes challenging in cluster-randomized experiments. This challenge is overcome in our proposed methodology, which is a conditional permutation test. We discuss this in Section 5.2.
@@ 5.1 Random graph models
This section describes various random graph models based on which $P(G)$ can be modeled. The Erdős-Rényi random graph model has been studied widely, owing to exact computations of desirable graph properties. We present the model formally in Supplementary material C.1. However, the Erdős-Rényi random graph model does not represent most real-world networks, such as social networks, supply-chain networks, and scientific collaboration networks, among many others. We look at a generalization of Erdős-Rényi random graph better at modeling such real-world networks, proposed by Newman et al. (2002), in Section 5.2.
5.2 Random graphs with arbitrary degree distribution
@@
Consider a random graph $G\in\{0,1\}^{N\times N}$ such that $P(G)=f(deg(i):i\in\mathbb{P}|p_{1},p_{2},...,p_{N})$. Here, $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution, that is, $p_{i}$ is the probability that a randomly selected unit has degree $i$ and $f(.)$ represents the probability mass function. We call $G$ a random graph with an arbitrary degree distribution.
A key feature of random graphs with arbitrary degree distribution is their ability to capture a large class of degree distributions. This generalizes the Erdős-Rényi random graph model and, at the same time, ensures tractability of important structural graph properties (Newman et al. (2002)). To deploy Procedure 1, we must identify $P(G)$. This involves estimating $(p_{1},p_{2},...,p_{N})$ with a single realization of the graph, and we cannot directly implement Procedure 1 without estimating $P(G)$. In the case of random graphs with arbitrary degree distribution, it is proven that we can estimate the parameters of the random graph with just one realization, but with some uncertainty (Chatterjee et al. (2011)). However, there might be other relational dependencies characterizing the formation of the network that are not captured by random graphs with arbitrary degree distribution. For example, the phenomena of transitivity are not captured in this model, where two nodes have higher chances of being connected if they have a common connecting node.
5.3 Exponential random graph model
@@
Common sufficient statistics used in the literature for the characterization of social networks include the number of edges, the number of triangles, and the number of stars of different sizes, say stars of size two or stars of size three (Robins et al. (2007)). It is important to carefully choose the model as an ill-posed ERGM can lead to near degeneracy (Chatterjee and Diaconis (2013); Schweinberger et al. (2020)). As with random graphs with arbitrary degree distribution, estimating ERGM is fraught with challenges, given just a single realization.
Owing to estimation issues of the null distribution, we propose a new method that does not involve such estimations and reduces to a simple permutation test. In our proposed methodology, we first consider the sufficient statistics of the random graph distribution as the degree sequence and develop a method conditioning on the sufficient statistic. We then generalize the method to a set of arbitrary sufficient statistics. We use the symmetric nature of the $P(G)$ via an equivalence relation and build a conditional permutation test, as described below.
6 Permutation test based on equiprobability events
6.1 Test for a random graph with arbitrary degree distribution
Here, we present our main result, which proposes obtaining a conditional randomization test without estimating the distribution. Our main idea is to identify a sufficient statistic for the random graph model, which is the degree sequence in the case of random graphs with arbitrary degree distribution, and condition on this sufficient statistic. We sample the null distribution over all the graphs with the same degree sequence as the observed graph. We obtain a permutation test using the equiprobability of the event set of all the graphs with the same degree sequence. We now formally present the definitions below and proceed to state and prove the finite sample validity of the proposed method.
@@ 6.1 Test for a random graph with arbitrary degree distribution
Not all $D\in\mathbb{W}^{N}$ can be a valid degree sequence. For example, consider $D=\{1,2\}$ as a candidate for a degree sequence for a graph of size $2$. We remark that the maximum degree of a node in the graph is at most (N1)(N-1). Since no vertex in a graph of size two can have a degree of more than 1, $D$ is not a valid degree sequence. Note that graphs with the same degree sequence are not necessarily isomorphic graphs. This implies that the graph’s vertices cannot be permuted to obtain the other graph with edges preserved (see Figure 1).
Figure 1: Non-isomorphic graphs with the same degree sequence
@@
Consider a random graph $G\in\{0,1\}^{N}\sim P(G)$ where $P(G)=f(deg(i):i\in V|p_{1},p_{2},...,p_{N})$ where $\{p_{i}\}_{i=1,...,N}$ represent the degree distribution of $G$. Let $\mathcal{G}$ be the support of $P(G)$. We define a relation $d:\mathcal{G}\rightarrow\mathcal{G}$ such that $G\equiv_{d}G^{\prime}$ if $deg(G)=deg(G^{\prime})$ such that $deg_{i}(G)=deg_{i}(G^{\prime})\forall i\in[N]$.
@@
    Define the equivalence class $G_{obs}]=\{G\in\mathcal{G}|\quad G\equiv_{d}G_{obs}\}$. Here, $\equiv_{d}$ is the relation defined in Definition 6.
@@
Consider the null hypothesis of no spillover effects in Hypothesis 4, denoted by $\mathcal{H}_{0}$. Let $(Z_{obs},G_{obs})\sim P(Z,G)$ be the bivariate treatment assignment of a completely randomized experiment. We assume that $G_{obs}\sim P_{G}$ where $G$ is a random graph with arbitrary degree distribution. Let $G_{X_{i}}=deg(i)$ where $G_{X_{i}}$ denote the graph covariate(s) of unit $i$. Consider a test statistic $T(Z,G,Y(Z,G))$. Then, Procedure 2 described above is conditionally valid at level $\alpha\in(0,1)$. That is,
@@
Refer to Supplementary material B.4. ∎
As a corollary, we can substitute any subclass of the degree sequence equivalence class in Procedure 2, and the finite-sample validity of the Procedure is retained. For example, we can take the equivalence class to be all the graphs isomorphic to the observed graph, or automorphic to the observed graph in the second step of Procedure 2. We present results for both these Procedures in Section 7.
We emphasize that generating graphs with a given degree sequence admits a polynomial-time algorithm. Given fixed parameters like the maximum degree of the graph, one can use state-of-the-art algorithms to efficiently sample graphs with a given degree sequence uniformly from the space of all the graphs with the same degree sequence (Arman et al. (2021)). It has been shown that the size of this equivalence class can be exponentially large (Barvinok and Hartigan (2013)). It should be noted that one should use a uniform sampler to avoid introducing any bias in the null distribution attributed to sampling. We can also overcome this by effectively taking many samples to approximate the null distribution.
6.2 Test for exponential random graph model
We generalize the above method to a class of exponential random graph models where the sufficient statistics of the model are permutation invariant. We can take the graph covariate vector, $G_{X}$, to be the same set of statistics. To sample graphs from $G|G_{X}$, we can condition on the graph automorphism class and generate the null cases from this class. One caveat is that the size of an automorphic graph class may not be large depending on the symmetry of the graph observed (Erdos and Rényi (1963)). To overcome this, we explore sampling approximate automorphic graphs. We fix the graph covariate to be the degree of a unit, and sample isomorphic graphs that preserve the degree sequence for the null cases. To achieve more generality, one can specify the graph covariate vector of interest and explore advanced graph sampling or rejection sampling techniques. We formally present the results below.